# Machine Learning Insights into Quark-Antiquark Interactions: Probing Field Distributions and String Tension in QCD

### v2

@@ I Introduction
Advances in artificial intelligence (AI), particularly in machine learning and deep learning, have benefited numerous fields in recent years, including physics. Neural networks have become a well-established tool for analyzing high-energy physics data [25], while deep learning is increasingly applied in high-energy phenomenology [26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Deep learning enhances flexibility in modeling data not easily described by explicit functions, even though its mechanisms often resemble a “black box” and remain difficult to fully interpret. Traditionally, model development for explaining experimental data or lattice QCD results has been theory-driven. In contrast, machine learning methods allow for the deduction of physical model features from data, constituting an “inverse problem” strategy. This data-driven approach reduces the risk of overly strong assumptions in model construction but still relies on human guidance for capturing physical characteristics. Neural networks based on multi-layer perceptrons (MLP) are commonly employed for parameter estimation, akin to global data fitting. Recently, the Kolmogorov-Arnold networks (KAN) method was introduced to the open-source community [36, 37], offering advantages in interpretability over MLP for small-scale scientific tasks and allowing human intervention during training. This study examines the similarities and differences between these two neural network methods in describing the flux tube profile between quark-antiquark pairs, with particular attention to how analytical expressions capture the distance dependence of the chromo field distribution. The paper is organized as follows: Sec. II reviews the chromo field distribution and introduces relevant aspects of machine learning. Sec. III presents the main machine learning results and numerical computations. The conclusion discusses findings and future directions.
@@ II.2 MLP vs KAN
In the MLP architecture of this work, the number of hidden layers is set to 2, each containing 128 neurons, and the activation function used is ReLU. This architecture is relatively straightforward to understand, and hyperparameters such as regularization, loss function, and learning rate can be adjusted according to the specific requirements of the training dataset.
@@ III.1 Feasibility assessment of the two machine learning architectures
The implementation of the MLP framework is illustrated in Fig. 1 (note that the input layer consists of a single neuron), where a standard training process can be readily found within the machine learning community. In contrast, the KAN architecture includes an additional step, specifically the determination of a symbolic representation of the objective function. For testing, the neuron structure of KAN is configured as $(1,3,3,1)$, as shown in Fig. 2. After initialization, each node can map to a spline function, which is the learnable component. Following one round of training, it is evident that only a single function path remains, a result of the penalty mechanism employed to eliminate high-frequency oscillations and highly irregular paths. Once a unique function path is obtained, pruning and refinement training can be performed, and upon reaching machine precision, linear regression can be conducted to determine the target symbolic function and its corresponding parameters. It is noteworthy that the loss functions corresponding to both the MLP and KAN frameworks in the univariate function training tests converge to $\sim 10^{-7}$.
@@ III.2 Two-dimensional input training of MLP and KAN
By comparing Eqs. (3) and (4), it is evident that both of them contain the univariate functions $\sin(x)$ and $\exp(x)$; however, Eq. (4) introduces another variable $d$ while omitting the quadratic function. Additionally, we observe that the training results of the KAN architecture with two-dimensional input are slightly inferior to those of the MLP architecture, as reflected in the convergence of the loss functions. Similar to the univariate MLP, the training loss function for the bivariate MLP converges to approximately $10^{-7}$, while in the case of KAN, the convergence is only around $10^{-2}$. We acknowledge that this discrepancy is strongly related to the hyperparameter dependence in the KAN training process. We also experimented with different hyperparameter combinations; one alternative option involves using a $(2,4,1)$ neuron configuration, which led to results that are independent of $d$. The training set itself exhibits limited dependence on $d$, as mentioned in the corresponding lattice studies [13, 15, 16], which is also reflected in the calculations of the flux tube string tension and the root mean square width based on the chromo field.
@@
$\sqrt{w^{2}}=\sqrt{\frac{\int d^{2}x_{t}x_{t}^{2}E_{(}x_{t})}{\int d^{2}x_{t}E_{(}x_{t})}}.$ (5)
@@ IV Summary and outlook
Understanding the spatial structure of gauge fields, particularly the flux tube model, is essential for elucidating the interactions between quark-antiquark pairs and the associated phenomena of quark confinement, including string breaking. In this study, we leverage machine learning techniques, notably MLP and the KAN, to analyze flux tube structures simulated via lattice QCD. Our objective is to harness AI to reproduce the lattice results effectively. We derive analytical expressions that characterize the chromo field distributions for varying quark-antiquark separations using the KAN method. We find that the MLP architecture adeptly captures the salient features of lattice data, while the KAN strategy involves greater complexity, particularly in replicating two-dimensional training datasets. Nonetheless, we successfully establish a two-dimensional expression for the chromo field distribution that closely mirrors lattice results. Additionally, we compute the variations of the flux tube string tension as a function of quark-antiquark separation, comparing outcomes from both trained MLP and KAN models. Our results reveal that the MLP approach outperforms KAN in operational efficiency and adaptability to large-scale datasets, while KAN offers advantages in generating interpretable analytical expressions with fewer assumptions. Despite the necessity for extensive optimization of KAN’s hyperparameters, we suggest the integration of adaptive hyperparameter tuning interfaces, such as Keras Tuner, to enhance the model’s applicability for learning physical phenomena. This work exemplifies the potential of AI methodologies in high-energy physics, contributing to the understanding of fundamental interactions via advanced computational techniques.

### v3

@@ I Introduction
Advances in artificial intelligence (AI), particularly in machine learning and deep learning, have benefited numerous fields in recent years, including physics. Neural networks have become a well-established tool for analyzing high-energy physics data [25], while deep learning is increasingly applied in high-energy phenomenology [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Deep learning enhances flexibility in modeling data not easily described by explicit functions, even though its mechanisms often resemble a “black box” and remain difficult to fully interpret. Traditionally, model development for explaining experimental data or lattice QCD results has been theory-driven. In contrast, machine learning methods allow for the deduction of physical model features from data, constituting an “inverse problem” strategy. This data-driven approach reduces the risk of overly strong assumptions in model construction but still relies on human guidance for capturing physical characteristics. Neural networks based on multi-layer perceptrons (MLP) are commonly employed for parameter estimation, akin to global data fitting. Recently, the Kolmogorov-Arnold networks (KAN) method was introduced to the open-source community [53, 54], offering advantages in interpretability over MLP for small-scale scientific tasks and allowing human intervention during training. This study examines the similarities and differences between these two neural network methods in describing the flux tube profile between quark-antiquark pairs, with particular attention to how analytical expressions capture the distance dependence of the chromo field distribution. In particular, the latter aspect is currently difficult to achieve with other parameterization schemes, which can only fit a set of parameters for specific quark-antiquark separations, resulting in poor continuity. Through deep learning, researchers can derive more continuous model representations from limited information or data. The paper is organized as follows: Sec. II reviews the chromo field distribution and introduces relevant aspects of machine learning. Sec. III presents the main machine learning results and numerical computations. The conclusion discusses findings and future directions.
@@ II.2 MLP vs KAN
In the MLP architecture of this work, the number of hidden layers is set to 2, each containing 128 neurons, and the activation function used is ReLU. This architecture is relatively straightforward to understand, and hyperparameters such as regularization, loss function, and learning rate can be adjusted according to the specific requirements of the training dataset. In this work, we configure the Mean Square Error (MSE) between the model predictions and lattice simulation results as the loss function, incorporating an L2 regularization scheme. To further validate the reliability of the training process and avoid overfitting,we randomly select $20\%$ of the data from the entire dataset as a test set for each training session. The trained model is required to describe both the test and training sets.
@@ III.1 Feasibility assessment of the two machine learning architectures
The implementation of the MLP framework is illustrated in Fig. 1 (note that the input layer consists of a single neuron), where a standard training process can be readily found within the machine learning community. In contrast, the KAN architecture includes an additional step, specifically the determination of a symbolic representation of the objective function. For testing, the neuron structure of KAN is configured as $(1,3,3,1)$, as shown in Fig. 2. After initialization, each node can map to a spline function, which is the learnable component. Following one round of training, it is evident that only a single function path remains, a result of the penalty mechanism employed to eliminate high-frequency oscillations and highly irregular paths. Once a unique function path is obtained, pruning and refinement training can be performed, and upon reaching machine precision, linear regression can be conducted to determine the target symbolic function and its corresponding parameters. It is noteworthy that the loss functions corresponding to both the MLP and KAN frameworks in the univariate function training tests converge to $\sim 10^{-5}$.
@@ III.2 Two-dimensional input training of MLP and KAN
By comparing Eqs. (3) and (4), it is evident that both of them contain the univariate functions $\sin(x)$ and $\exp(x)$; however, Eq. (4) introduces another variable $d$ while omitting the quadratic function. Additionally, we observe that the training results of the KAN architecture with two-dimensional input are slightly inferior to those of the MLP architecture, as reflected in the convergence of the loss functions. Similar to the univariate MLP, the training and testloss functions for the bivariate MLP both converges to approximately $10^{-5}$ (see Fig. 5), while in the case of KAN, the convergence is only around $10^{-2}$. We acknowledge that this discrepancy is strongly related to the hyperparameter dependence in the KAN training process. We also experimented with different hyperparameter combinations; one alternative option involves using a $(2,4,1)$ neuron configuration, which led to results that are independent of $d$. The training set itself exhibits limited dependence on $d$, as mentioned in the corresponding lattice studies [13, 15, 16], which is also reflected in the calculations of the flux tube string tension and the root mean square width based on the chromo field.
Figure 5: The comparison of the loss functions between the training and test sets in the case of MLP architecture.
@@
$\sqrt{w^{2}}=\sqrt{\frac{\int d^{2}x_{t}x_{t}^{2}E(x_{t})}{\int d^{2}x_{t}E(x_{t})}}.$ (5)
@@ IV Summary and outlook
Understanding the spatial structure of gauge fields, particularly the flux tube model, is essential for elucidating the interactions between quark-antiquark pairs and the associated phenomena of quark confinement, including string breaking. In this study, we leverage machine learning techniques, notably MLP and the KAN, to analyze flux tube structures simulated via lattice QCD. Our objective is to harness AI to reproduce the lattice results effectively. We derive analytical expressions that characterize the chromo field distributions for varying quark-antiquark separations using the KAN method. We find that the MLP architecture adeptly captures the salient features of lattice data, while the KAN strategy involves greater complexity, particularly in replicating two-dimensional training datasets. Nonetheless, we successfully establish a two-dimensional expression for the chromo field distribution that closely mirrors lattice results. Previously, the parameterization of chromo fields based on the second type of superconducting duality could only provide a set of parameters corresponding to a fixed quark-antiquark separation $d$. In other words, while these models describe lattice results well, they are limited in predicting the chromo field profiles for unknown values of $d$. We believe that applying a neural network deep learning approach to analyze the existing limited lattice simulations can contribute positively to predicting the chromo field distributions for different values of $d$.
Additionally, we compute the variations of the flux tube string tension as a function of quark-antiquark separation, comparing outcomes from both trained MLP and KAN models. Our results reveal that the MLP approach outperforms KAN in operational efficiency and adaptability to large-scale datasets, while KAN offers advantages in generating interpretable analytical expressions with fewer assumptions. Despite the necessity for extensive optimization of KAN’s hyperparameters, we suggest the integration of adaptive hyperparameter tuning interfaces, such as Keras Tuner, to enhance the model’s applicability for learning physical phenomena. This work exemplifies the potential of AI methodologies in high-energy physics, contributing to the understanding of fundamental interactions via advanced computational techniques.